Project

General

Profile

11192013 Minutes

Attending: Micheal, Dom, Nick, Evan, Siva, Mat, Kanika, Gav, Alex

Operations:
-Noisy detector is returning back to normal.
-We are returning to running with one partition.
-DCM 04 on Diblock
-Siva: Will work with Gavin today

Production: No reco job for the last 40 runs.
- FTS is work well for last 3 days.

Event scanning: CAF ana macros, with event filtering (before)
: Now there is an Art Filter module
: Tutorial tomorrow.

Evan: Approval for 2 new keytab files. nova_mon_xfr. Scienfic Computing approved.
- Committed a fix to the bad change service to load a run specific.
- Data bad channel mask look like it’s working. 11600’s and 11500’s had different masks.

Micheal - Micheal has been manually updating onmon.